OpenStudy (anonymous):

How do I solve lnx+ ln(2x+1)=0

OpenStudy (anonymous):

the answer is 1/2 but I don't know how to get there

jimthompson5910 (jim_thompson5910):

ln(x)+ ln(2x+1)=0 ln(x*(2x+1))=0 x*(2x+1) = e^0 x*(2x+1) = 1 2x^2 + x = 1 2x^2 + x - 1 = 0 I'll let you finish. Don't forget to check your answers.

OpenStudy (anonymous):

I solved, got the answer, and understand everything except the RHS of the third step...how do I know that 0 is supposed to be e^0 and that e^0 means one? or what does e mean?

jimthompson5910 (jim_thompson5910):

I'm using the idea that if ln(x) = y then x = e^y

jimthompson5910 (jim_thompson5910):

since in general \[\large \log_{b}(x) = y\] turns into \[\large x = b^y\]

OpenStudy (jdoe0001):

\(\bf ln(x \times (2x+1))=0 \implies log_e(x \times (2x+1))=0\\ \textit{exponentialize both sides by "e"}\\ \color{blue}{\Large e^{log_e(x \times (2x+1))} = e^{0}}\\ \textit{log cancellation rule of } \Large a^{log_ax} = x\\ x \times (2x+1) = e^{0}\)

OpenStudy (anonymous):

Thanks to both of you!

jimthompson5910 (jim_thompson5910):

yw